vector perpendicular to 4i-3j is ??

Grade:

3 Answers

Chetan Mandayam Nayakar
11 years ago

let answer=ai+bj

4a-3b=0

nikkitotti
11 years ago

(4i-3j). vector a=0

a=ai+bj

(4i-3j).(ai+bj)=0

4a-3b=0

a=3b/4

ans- 3b/4i+bj

 

Rohit Gupta
11 years ago

Then Vector pependicular to this is z axis;

i.e. (--)k.